OpenStudy (anonymous):

SHOW ALL OF THE STEP THAT YOU USE TO SOLVE THE PROBLEM 20. The midpoint of UV is (5-11). The coordinates of one endpoint are U(3,5). FIND the coordinates of endpoint V.

OpenStudy (anonymous):

\((3,5)\) is the endpoint, \((5,-11)\) is the midpoint?

OpenStudy (anonymous):

if so , you can almost do it in your head from \(3\) to \(5\) in the \(x\) direction is right two units if you go another two units to the right from \(5\) you get to \(7\)

OpenStudy (anonymous):

from \(5\) to \(-11\) in the \(y\) direction is down \(-16\) units go down another 16 and get to \(-27\) the other endpoint is therefore \((7,-27)\)

OpenStudy (anonymous):

unless of course the midpoint was \((5,11)\) in which case the answer is different, not really clear from the question

OpenStudy (anonymous):

is it \((3,5)\) and \((5,11)\) or is it \((3,5)\) and \((5,-11)\) ?

OpenStudy (chris911):

long story short is there a picture or steps they give u ??

OpenStudy (anonymous):

noo

OpenStudy (chris911):

hang on a sec

OpenStudy (chris911):

try this Since the midpoint is (5, –11), it implies the total length is 2(5, -11) = (10, -22) But U(3, 5). This implies V(x, y) = (10, -22) + (3, 5) = (13, -17) Therefore the coordinate of V is (13, -17)

OpenStudy (anonymous):

Let the other end point be x2, y2 : Find x2 3+x2/2 = 5 Multiply both sides by 2 3 + x2 = 10 x2 = 10 - 3 x2 = 7 : Find y2 5+y2/2 = -11 multiply both sides by 2 5 + y2 = -22 y2 = -22 - 5 y2 = -27 : V endpoint: 7, -27

OpenStudy (chris911):

or this we have the end point at (0,0) then the midpoint is easy to work with :) (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) since the midpoint is half way, double it to get to the end. (3, 5) (5,–11) (x , y ) -3 -5 -3 -5 -3 -5 ----------------------- (0, 0) (2,-16) (x-3, y-5) 4 -32 --------- (x-1 , y-37) we modified (x=3,y=5) to begin with so lets sub those values back in. (3-1, 5-37) = (2, -27)

OpenStudy (anonymous):

thx you

OpenStudy (anonymous):

wow that is a lot of work!

OpenStudy (anonymous):

solve \(\frac{3+x}{2}=5\) get \(3+x=10\) or \(x=7\) for the first coordinate

OpenStudy (anonymous):

for the second one, solve \[\frac{y+5}{2}=-11\] get \(y+5=-22\) and so \(y=27\) your answer is \((7,-27)\) and that is all